OpenStudy (anonymous):

Can someone help me out please? im working on Graphing Exponential Functions...

OpenStudy (jessicawade):

what is the question

OpenStudy (solomonzelman):

It will look like a curve, or if it a quadratic equation it will be a parabola that is opening up or down.

OpenStudy (anonymous):

Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days. 1.Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. 2.Graph the price function for Stock D. Explain the key features of the graph with complete sentences. 3.A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). 4.Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. •f(x + 2) •f(–x) •f(x) + 3

OpenStudy (anonymous):

@jessicawade @SolomonZelman

OpenStudy (solomonzelman):

Stock D ?

OpenStudy (anonymous):

The other section on the image I sent where the columns are labled abc @SolomonZelman

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ \rm Stock~~A:~~~~f(x)=25(1.08)^x}\) 8% increase. \(\normalsize\color{black}{ \rm Stock~~B:~~~~f(x)=25(0.98)^x}\) ≈2% decrease. \(\normalsize\color{black}{ \rm Stock~~C:~~~~f(x)=30(1.04)^x}\) 4% increase.

OpenStudy (solomonzelman):

I don't see stock D -:(

OpenStudy (anonymous):

I have to make stock d The instructions are in question1. 1.Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Okya

OpenStudy (solomonzelman):

second lower starting price . btw, \(\normalsize\color{black}{ \rm Stock~~B:~~~~f(x)=22(0.98)^x}\)

OpenStudy (solomonzelman):

So rvwing \(\normalsize\color{black}{ \rm Stock~~A:~~~~f(x)=25(1.08)^x}\) x=0, f(x)=25 \(\normalsize\color{black}{ \rm Stock~~B:~~~~f(x)=22(0.98)^x}\) x=0, f(x)=22 \(\normalsize\color{black}{ \rm Stock~~C:~~~~f(x)=30(1.04)^x}\) x=0, f(x)=30

OpenStudy (solomonzelman):

(When x=0, that is the starting price )

OpenStudy (solomonzelman):

x=0 is the starting time, and f(X) is the price at a certain time "x"

OpenStudy (solomonzelman):

So your initial price has to be between 22 and 25. Will can this price 'v'. And 22<v<25. (or same thing as in an interval (22,25) )

OpenStudy (solomonzelman):

And second highest growing rate, so it has to be between the rate of of Stock B and the grown rate of Stock C.

OpenStudy (solomonzelman):

It has to be more than -2%, and less than +4%.

OpenStudy (solomonzelman):

We can say, 0.98 < common ratio < 1.04 Makes sense ?

OpenStudy (solomonzelman):

bye, have fun ... !

OpenStudy (anonymous):

Sorry my computer froze.. yes thank you! @SolomonZelman